1. An unknown noble gas has a density of 5.84 g dm–3 at STP. Calculate its molar mass, and so identify the gas. 
1
Expert's answer
2021-02-22T05:50:45-0500

Density of unknown gas("\rho") = 5.84 g dm-3 = 5.84 g/L

Let molar mass of unknown gas = M g/mol

At standard temperature and pressure conditions,

Temperature(T) = 273.15 K

Pressure(P) = 1 atm

Gas Constant(R) = 0.0821 L atm K-1 mol-1

From ideal gas equation,

"\Rightarrow PV= nRT"

"\Rightarrow PV= \dfrac{m}{M}RT"


"\Rightarrow M= \dfrac{m}{V}\dfrac{RT}{P}"


"\Rightarrow M = \rho \dfrac{RT}{P}"

"\Rightarrow M = 5.84 \times \dfrac {0.0821 \times 273.15}{1}"

"\Rightarrow M = 130.965g/mol \approx 131 g/mol"


Xenon has molar mass approximately equal to 131 g/mol